What will be the final temperature if 50g of water at exactly 0°c is added to 250g of water at 90°c?

Grade:10

2 Answers

Mohamed Harithkhan
4 years ago
If heat lose/gain from the environment,
 
Heat lost by  250g of water at 90°c will be equal to heat gained by  50g of water at exactly 0°c.
Let’s take final temperature as t
So, m1*S*Δt1 = m2*S*Δt2
250 * (90 – t) = 50 * (t – 0)
300t = 90*250
t = 75°c
So, final temperature is 75°c
Mohamed Harithkhan
4 years ago
If heat lose/gain from the environment are neglected,
 
Heat lost by  250g of water at 90°c will be equal to heat gained by  50g of water at exactly 0°c.
Let’s take final temperature as t
So, m1*S*Δt= m2*S*Δt2
250 * (90 – t) = 50 * (t – 0)
300t = 90*250
t = 75°c
So, final temperature is 75°c